TIFF2BW(1) TIFF2BW(1) NAME tiff2bw−convertacolorTIFFimagetogreyscale SYNOPSIS tiff2bw[options]input.tifoutput.tif DESCRIPTION Tiff2bwconvertsanRGBorPalettecolorTIFFimagetoagreyscaleimagebycombiningpercentagesofthered,green,andbluechannels.Bydefault,outputsamplesarecreatedbytaking28% oftheredchannel,59%ofthegreenchannel,and11%ofthebluechannel.Toalterthesepercentages,the−R,−G,and−Boptionsmaybeused. OPTIONS −cSpecifyacompressionschemetousewhenwritingimagedata:−cnonefornocompression,-cpackbitsforthePackBitscompressionalgorithm,-czipfortheDeflatecom-pressionalgorithm,-cg3fortheCCITTGroup3compressionalgorithm,-cg4forthe CCITTGroup4compressionalgorithm,and−clzwforLempel-Ziv&Welch(the default). −rWritedatawithaspecifiednumberofrowsperstrip;bydefaultthenumberofrows/strip isselectedsothateachstripisapproximately8kilobytes. −RSpecifythepercentageoftheredchanneltouse(default28). −GSpecifythepercentageofthegreenchanneltouse(default59). −BSpecifythepercentageofthebluechanneltouse(default11). SEEALSO pal2rgb(1),tiffinfo(1),tiffcp(1),tiffmedian(1),libtiff(3) October15,1995 1